OpenStudy (anonymous):

Rationalize the Denominator sqrt(14/3)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Oh and an explination would be nice.

OpenStudy (amistre64):

\[\sqrt{\frac{14}{3}}\] this expression?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

the radical splits to this:\[\frac{\sqrt{14}}{\sqrt{3}}\]

OpenStudy (anonymous):

I understood that far.. Didn't know where to go from there.

OpenStudy (amistre64):

then to get rid of the \(\sqrt{(..)}\) in th bottom; we multiply top and bottom by it

OpenStudy (amistre64):

\[\frac{\sqrt{14}}{\sqrt{3}}*\frac{\sqrt{3}}{\sqrt{3}}=\frac{\sqrt{14*3}}{3}\]

OpenStudy (anonymous):

Thanks man. That makes sense now.

OpenStudy (amistre64):

\[\sqrt{14*3}=\sqrt{42}\] whih may or maynot simplify more :)

OpenStudy (anonymous):

sqrt(42) is simplest

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[\sqrt{42}\] is the simplest because 42 is a multiplication of 2 prime number :14 and 3. Amistre provided clear steps. :)

OpenStudy (anonymous):

14 is not a prime no as it can be split as 2x7 !!!!!

OpenStudy (anonymous):

@amistre after rationalising where has the 3 from the denominator vanished????

OpenStudy (anonymous):

final answer shud b root42 ------- 3